OpenStudy (anonymous):

given triangle ABC whose sides are AB=15in., AC=25in., and BC=30in., . from a point D on side AB, a line DE is drawn to a point E on side AC such that angle ADE is equal to angle ABC. if the perimeter of triangle ADE is 28m. .find the lengths of line segments BD and CE

OpenStudy (amistre64):

sounds like similar triangles to me

OpenStudy (amistre64):

as such, the perimeter of one is just a scaled version of the other

OpenStudy (anonymous):

i can sketch it :(

OpenStudy (amistre64):

draw a triangle with a line parallel to the stated side

OpenStudy (amistre64):



OpenStudy (anonymous):

we are the same

OpenStudy (amistre64):

good, as such we are looking to find a scalar value to find AD and AE to finish the solution

OpenStudy (anonymous):

the ADE is form a scalene triangle

OpenStudy (amistre64):

the scaled parts are such that for some value k: 15k + 25k +30k = 28 ... is that meters? if so we would want to convert it to inches so be in the same dimension as the stated side values

OpenStudy (amistre64):

check to make sure thats spose to be 28m, and not 28in otherwise check the stated side values and see if their units values are correct

OpenStudy (anonymous):

yes sir thats inches

OpenStudy (amistre64):

so 28 inches is a better situation to play with ... no conversions needed :) so lets solve for k

OpenStudy (amistre64):

15k + 25k +30k = 28 k(15 + 25 +30) = 28 k(70) = 28 .... whats the value for k?

OpenStudy (anonymous):

2/5

OpenStudy (amistre64):

2/5 i like ... soo now the question is reduced to: BA = BD + AD, but AD = BAk BA = BD + BAk BA - BAk= BD, is one answer, same concept applies to the other side

OpenStudy (anonymous):

9 and 15???

OpenStudy (amistre64):

9 is good: 15-15*2/5 = 15-6 = 9 25-25*2/5 = 25-10 = 15, i like it

OpenStudy (anonymous):

i got it sir thank you :)

OpenStudy (amistre64):

youre welcome, and good luck :)

OpenStudy (anonymous):

your the best :)